a quart of vodka is about 40% alcohol. how much vodka should be mixed with orange juice to make 2 gallons of screwdriver mix with an alcohol content of 10%
Answer by stanbon(75887)

a quart of vodka is about 40% alcohol. how much vodka should be mixed with orange juice to make 2 gallons of screwdriver mix with an alcohol content of 10%
--------------------
Equation:
alcohol + alcohol = alcohol
0.40x + 0(2-x) = 0.10*2
0.40x = 0.20
x = (1/2) gallons (amt. of vodka needed for the mixture)
==============================================================
Cheers,
Stan H.